In a class of 30 students, the number of students who can speak English is 16, and the number of those who speak French is 5. Apart from these, there are no students who can speak a foreign language. In this class:

a) What is the least probability of those who do not speak a foreign language?

b) What is the maximum probability of people who do not speak a foreign language?

c) What is the maximum probability of those who know both foreign languages?

 d) What is the minimum probability of people who know only one foreign language?

e) What is the maximum probability of people who speak only one foreign language?


1
Expert's answer
2021-01-12T12:31:59-0500

Total strength = 30

No of English speakers = 16

No of French speakers = 5

a) Here as per the case we have to consider that all English speakers cant speak French and all French speaker cant speak English

Required probability "= \frac{30-(16+5)}{30}"

"= \frac{3}{10} \\

= 0.30"

b) Here as per the case we have to consider that all all French speakers can speak English too

Therefore

No of non foreign language speakers = 30-16 = 14

Required probability "= \frac{14}{30} = 0.466"

c) Since there are only 5 people who can speak French. So there can be at max 5 people who can speak both French and English

Required probability "= \frac{5}{30}= 0.166"

d) In this case we consider that all French speakers cant speak English

Therefore

Required probability "= \frac{5}{30}= 0.166"

e) In this case we will consider that all English speakers don't know to speak French

Therefore

Required probability "= \frac{16}{30}= 0.533"